46. find the values of sinu/2 cosu/2 tanu/2
cosu=3/5

Answer by stanbon(57246)

find the values of [sin(u) / 2cos(u)] / [2tan(u) / 2cos(u)] =3/5
-----------------
Invert the denominator and multiply:
[sin(u) / 2cos(u)] * [2cos(u)/2tan(u)] =3/5
Simplify:
[(1/2)tan(u)] * [2cos(u)/2tan(u)] =3/5
Cancel the common factors of 2 and tan(u) to gt:
------
[(1/2)] * [cos(u)/1] = 3/5
cos(u) = 6/5
------------------
But cos values must be between -1 and +1.
Conclusion: No solution
---------------------
Cheers,
Stan H.